OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = -10 3x - 3y + 6z = -21 x + 2y + 2z = -10 (1, -2, -3) (-4, -1, -2) (-3, -2, -3) (2, -2, -4)

OpenStudy (zzr0ck3r):

linear algebra?

OpenStudy (zzr0ck3r):

you can make a matrix and reduce it, but I would just check the answers to see which once works......

OpenStudy (mathmate):

Add all equations to give 5x+3y+9z=-41 A: 5(1)+3(-2)+9(-3)=-28 not the right one. B: 5(-4)+3(-1)+9(-2)=? C: 5(-3)+3(-2)+9(-3)=? D: 5(2)+3(-2)+9(-4)=? Continue with ALL the other choices. There may be more than one that fits the bill. If that happens, check each individually.